OpenStudy (anonymous):

please just check to see if i got the right answer! a tangent drawn to y=-(x^2 -4) at the point (1, 3) forms a right triangle with the coordinate axes. find the area of this right triangle. please help!

OpenStudy (mathmate):

What is your answer?

OpenStudy (anonymous):

oh sorry i said that the area is 16

OpenStudy (mathmate):

f(x)=-(x^2-4) f'(x)=-2x at (1,3), the tangent line is: y=-2(x-1)+3=-2x+5 the x-intercept is 5/2, and the y-intercept is 5 So the area is (5/2)(5) / 2 = 25/4